# competing interactions; square well; square shoulder; discontinuous structural crossover transitions; Fisher–Widom line; Rational-Function Approximation

### v1

@@ 3.1 Theoretical background
Figure 5(a) displays the DOC lines for 3D fluids with $\epsilon_{1}=\epsilon_{2}>0$, corresponding to values of $\lambda_{2}=1.55$, $1.6$, $1.65$, $1.7$, $1.75$, and $1.8$. The overall shape of these lines is qualitatively similar to the lines shown in Figure 1(b) for $\lambda_{2}=1.8$ and $1.7$, but with noticeably smaller loops, particularly as $\lambda_{2}$ increases. Within these loops, as in the 1D case, the oscillation frequency is approximately $\omega/2\pi\approx 3/\lambda_{2}$. Furthermore, the DOC arcs observed in Figures 1(b) and 1(c) for 1D fluids are absent in Figure 5(a), as they would be confined to the high-density, low-temperature region where the RFA is no longer reliable. Indeed, no DOC line was observed for $\lambda_{2}\leq 1.5$, consistent with the disappearance of the single DOC line in Figure 1(c) for $\lambda_{2}=1.6$ and $1.5$. Additionally, the 3D density playing the role of the 1D value $\rho^{*}=\lambda_{2}^{-1}$ would be given by $\rho^{*}=\rho^{*}_{\max}\lambda_{2}^{-3}$, where $\rho^{*}_{\max}\simeq 0.94$ represents the freezing density of hard spheres Alder and Wainwright (1957); Robles et al. (2014).
@@
A comparison between Figure 5(b) and the inset of Figure 1(a) reveals a shared characteristic: when the single DOC line is crossed at a given density while moving from higher to lower temperatures, the frequency $\omega$ initially increases near the crossover temperature before suddenly dropping to a smaller value.
@@
Figure 5: (a) DOC lines on the $T^{*}$ vs $\rho^{*}$ plane for the 3D case $\epsilon_{1}=\epsilon_{2}$ with $\lambda_{2}=1.55$, $1.6$, $1.65$, $1.7$, $1.75$, and $1.8$. The inset shows the loop corresponding to $\lambda_{2}=1.55$. (b) Angular frequency of the asymptotic oscillations of $h(r)$ plotted as a function of $T^{*}$ for $\rho^{*}=0.05$, $0.1$, $0.15$, $0.2$, $0.25$, $0.3$, and $0.4$, with an interaction potential characterized by $\epsilon_{1}=\epsilon_{2}$ and $\lambda_{2}=1.7$. The arrows indicate the direction of increasing (a) $\lambda_{2}$ and (b) $\rho^{*}$.
3.3 $\lambda_{1}=1.35$, $\lambda_{2}=1.7$. Influence of $\epsilon_{1}/\epsilon_{2}$ on the DOC line
@@
Figure 6(a) displays the DOC lines on the $T^{*}$ vs $\rho^{*}$ plane for various values of $\epsilon_{1}/\epsilon_{2}$, covering the cases where $\epsilon_{1}>0$, $\epsilon_{1}=0$, and $\epsilon_{1}<0$. In analogy with the 1D case [see Figures 3(b) and 3(c)], these lines exhibit qualitative changes as the system transitions from positive to negative values of $\epsilon_{1}$. However, in the 3D case, apparently the loops do not degenerate into lobes.
@@
Figure 6: (a) DOC lines on the $T^{*}$ vs $\rho^{*}$ plane for the 3D case $\lambda_{1}=1.35$ and $\lambda_{2}=1.7$ with, from right to left, $\epsilon_{1}/\epsilon_{2}=2.5$, $2$, $1$, $0.5$, 00, $-0.5$, $-0.75$, $-1$, and $-2$. The inset shows the loop corresponding to $\epsilon_{1}/\epsilon_{2}=-0.75$. (b) Angular frequency of the asymptotic oscillations of $h(r)$ plotted as a function of $T^{*}$ for $\rho^{*}=0.04$, $0.13$, and $0.14$, with an interaction potential characterized by $\epsilon_{1}/\epsilon_{2}=-0.75$, $\lambda_{1}=1.35$, and $\lambda_{2}=1.7$.
@@
Another notable feature is the rounded, bulging profile of the DOC lines for $\epsilon_{1}<0$. This shape indicates that, within a certain density interval, the frequency $\omega$ exhibits reentrant behavior as the temperature varies. This phenomenon is illustrated in Figure 6(b) for $\epsilon_{1}/\epsilon_{2}=-0.75$. At a density $\rho^{*}=0.04$ (below the loop densities $\rho^{*}\approx 0.05$), the oscillation frequency undergoes a single drop from $\omega\simeq 6.6$ to $\omega\simeq 2.1$ when crossing the temperature $T^{*}\simeq 0.76$ from left to right, and then increases smoothly toward $\omega\simeq 4.4$ in the high-temperature limit. At a higher density, $\rho^{*}=0.13$ (just below the bulge’s end at $\rho^{*}\simeq 0.138$), a more complex behavior is observed: the frequency drops from $\omega\simeq 6.6$ to $\omega\simeq 2.5$ at $T^{*}\simeq 1.3$, then rises again from $\omega\simeq 2.7$ to $\omega\simeq 6.5$ at $T^{*}\simeq 3.6$, eventually tending smoothly toward $\omega\simeq 4.9$ at high temperatures. Finally, at $\rho^{*}=0.14$, the evolution of $\omega$ from $\omega\simeq 6.6$ at low temperatures to $\omega\simeq 4.9$ at high temperatures proceeds continuously without reentrant behavior.
3.4 $\lambda_{1}=1.35$, $\lambda_{2}=1.7$, $\epsilon_{1}<0$. Influence of $\epsilon_{1}/\epsilon_{2}$ on the FW line
@@
In the 1D case, an FW line was already observed with $\epsilon_{1}/\epsilon_{2}=-0.5$, but this required temperatures on the order of $T^{*}\sim 10^{-1}$ [see Figure 4(a)]. Since, as mentioned earlier, the RFA tends to provide less reliable results at low temperatures, it becomes necessary to consider deeper wells to study the FW lines for 3D fluids. The cases $\epsilon_{1}/\epsilon_{2}=-4$, $-8$, $-20$, and $-50$ are reported in Figure 7(a). As in the 1D fluid, it is useful to plot the curves on the $T_{1}^{*}$ vs $\rho^{*}$ plane to compare them with the FW line of the pure square-well fluid, as shown in Figure 7(b). Again, we observe that the presence of the repulsive barrier between $\lambda_{1}$ and $\lambda_{2}$ bends the FW line downward in the low-density region. This indicates that the decay of $h(r)$ is always oscillatory when the temperature exceeds a certain threshold, regardless of the density.
@@
Figure 7: (a) FW lines on the $T^{*}$ vs $\rho^{*}$ plane for the 3D case $\lambda_{1}=1.35$ and $\lambda_{2}=1.7$ with, from bottom to top, $\epsilon_{1}/\epsilon_{2}=-4$, $-8$, $-20$, and $-50$. (b) Same as in panel (a), except that now the vertical axis represents the scaled temperature $T_{1}^{*}=k_{B}T/|\epsilon_{1}|=T^{*}\epsilon_{2}/|\epsilon_{1}|$. The dotted line is the FW line for a pure square-well fluid ($\epsilon_{2}/|\epsilon_{1}|=0$).
@@ 4 Conclusions
In this paper, we have explored the impact of competing interactions in the intermolecular potential of fluids on their structural transitions. The model potential adopted for both 1D and 3D systems consists of a hard core followed by two steps, which can represent either a shoulder or a well depending on the sign of the parameters $\epsilon_{1}$ and $\epsilon_{2}$. This potential is versatile enough to encompass a range of competing interactions, including the square-well and square-shoulder interactions as limiting cases. Additionally, the consideration of two different dimensionalities allowed us to examine the influence of strong confinement on the structural transitions of these fluids. For the 1D systems, restricting the interaction range to no more than twice the hard-core diameter enabled us to derive exact results. In contrast, for the 3D systems, where exact solutions are not feasible, we employed the RFA to obtain and analyze approximate structural properties.
@@
The results for both the 1D and 3D systems align with the expected behavior. Specifically, at very low temperatures, the decay of the total correlation function $h(r)$ exhibits oscillations with a wavelength determined by the range of the repulsive barrier, provided that both $\epsilon_{1}$ and $\epsilon_{2}$ are positive. In contrast, at very high temperatures, the oscillations have a wavelength on the order of the hard-core diameter. Furthermore, it is confirmed that at a given density, the transition between these two regimes as the temperature increases can occur either continuously or, as observed in binary hard-sphere mixtures, discontinuously upon crossing a DOC line.
@@
When $\epsilon_{1}$ is negative, an FW transition from an oscillatory to a monotonic decay of $h(r)$ occurs as the temperature decreases at a given density, even when $\epsilon_{2}$ is positive. Additionally, the presence of the repulsive barrier of height $\epsilon_{2}$ causes the FW line to exhibit a maximum at a certain density before bending downward at lower densities, in stark contrast to its behavior in the absence of such a barrier.
@@
While the results for both the 1D and 3D systems exhibit many common characteristic features, the effects of dimensionality introduce notable distinctions. These include shifts in the temperature ranges at which certain features appear, the need for deeper wells to observe similar phenomena, or a reduction in their prominence as the system transitions from 1D to 3D. In some cases, features present in 1D may vanish entirely in 3D. Notably, we emphasize the complex behavior of the DOC transition, as previously discussed. This intricacy manifests in phenomena such as loops, arcs, lobes, triple points, and reentrant frequencies—some of which, to the best of our knowledge, have not been reported in this context before.
@@
In summary, we have uncovered a remarkably complex pattern of structural transitions in fluids with intermolecular potentials that include competing interactions. Even for the relatively simple potential considered in this work, analyzing structural transitions requires exploring a broad (dimensionless) parameter space, involving $\lambda_{1}$, $\lambda_{2}$, $\epsilon_{1}/\epsilon_{2}$, $T^{*}$, and $\rho^{*}$. Given these circumstances, our findings should be viewed as an initial exploration—indicative but necessarily incomplete. Nonetheless, they highlight a fascinating and intricate phenomenology that warrants further and more detailed investigation.
@@
Conceptualization, A.S.; methodology, S.B.Y., A.S., and M.L.H.; software, A.M.M. and S.B.Y.; validation, A.M.M., S.B.Y., A.S., and M.L.H.; formal analysis, A.S. and M.L.H.; investigation, A.M.M., S.B.Y., A.S., and M.L.H.; writing—original draft preparation, A.S. and M.L.H; writing—review and editing, A.M.M., S.B.Y., A.S., and M.L.H.; visualization, A.M.M.; supervision, S.B.Y., A.S., and M.L.H.; funding acquisition, S.B.Y. and A.S. All authors have read and agreed to the published version of the manuscript.
@@
This research was funded by MCIN/AEI/10.13039/501100011033 grant number PID2020-112936GB-I00. A.M.M. is grateful to the Spanish Ministerio de Ciencia e Innovación for a predoctoral fellowship Grant No. PRE2021-097702.
@@
Appendix A Some mathematical details in the case of the 1D fluid.

### v2

@@ 3.2 $\epsilon_{1}=\epsilon_{2}>0$: Influence of $\lambda_{2}$ on DOC Line
Figure 5a displays the DOC lines for 3D fluids with $\epsilon_{1}=\epsilon_{2}>0$, corresponding to values of $\lambda_{2}=1.55$, $1.6$, $1.65$, $1.7$, $1.75$, and $1.8$. The overall shape of these lines is qualitatively similar to the lines shown in Figure 1b for $\lambda_{2}=1.8$ and $1.7$ but has noticeably smaller loops, particularly as $\lambda_{2}$ increases. Within these loops, as in the 1D case, the oscillation frequency is approximately $\omega/2\pi\approx 3/\lambda_{2}$. Furthermore, the DOC arcs observed in Figure 1b,c for 1D fluids are absent in Figure 5a, as they would be confined to the high-density, low-temperature region where the RFA is no longer reliable. Indeed, no DOC line is observed for $\lambda_{2}\leq 1.5$, consistent with the disappearance of the single DOC line in Figure 1c for $\lambda_{2}=1.6$ and $1.5$. Additionally, the 3D density playing the role of the 1D value $\rho^{*}=\lambda_{2}^{-1}$ is given by $\rho^{*}=\rho^{*}_{\max}\lambda_{2}^{-3}$, where $\rho^{*}_{\max}\simeq 0.94$ represents the freezing density of hard spheres Alder and Wainwright (1957).
@@
A comparison between Figure 5b and the inset of Figure 1a reveals a shared characteristic: when the single DOC line is crossed at a given density while moving from higher to lower temperatures, the frequency $\omega$ initially increases near the crossover temperature before suddenly dropping to a smaller value.
@@
Figure 5: (a) DOC lines on the $T^{*}$ vs. $\rho^{*}$ plane for the 3D case $\epsilon_{1}=\epsilon_{2}$ with $\lambda_{2}=1.55$, $1.6$, $1.65$, $1.7$, $1.75$, and $1.8$. The inset shows the loop corresponding to $\lambda_{2}=1.55$. (b) The angular frequency of the asymptotic oscillations of $h(r)$ plotted as a function of $T^{*}$ for $\rho^{*}=0.05$, $0.1$, $0.15$, $0.2$, $0.25$, $0.3$, and $0.4$, with an interaction potential characterized by $\epsilon_{1}=\epsilon_{2}$ and $\lambda_{2}=1.7$. The arrows indicate the direction of increasing (a) $\lambda_{2}$ and (b) $\rho^{*}$.
3.3 $\lambda_{1}=1.35$ and $\lambda_{2}=1.7$: Influence of $\epsilon_{1}/\epsilon_{2}$ on DOC Line
@@
Figure 6a displays the DOC lines on the $T^{*}$ vs. $\rho^{*}$ plane for various values of $\epsilon_{1}/\epsilon_{2}$, covering the cases where $\epsilon_{1}>0$, $\epsilon_{1}=0$, and $\epsilon_{1}<0$. In analogy with the 1D case [see Figure 3b,c], these lines exhibit qualitative changes as the system transitions from positive to negative values of $\epsilon_{1}$. However, in the 3D case, the loops apparently do not degenerate into lobes.
@@
Figure 6: (a) DOC lines on the $T^{*}$ vs. $\rho^{*}$ plane for the 3D case $\lambda_{1}=1.35$ and $\lambda_{2}=1.7$ with, from right to left, $\epsilon_{1}/\epsilon_{2}=2.5$, $2$, $1$, $0.5$, 00, $-0.5$, $-0.75$, $-1$, and $-2$. The inset shows the loop corresponding to $\epsilon_{1}/\epsilon_{2}=-0.75$. (b) The angular frequency of the asymptotic oscillations of $h(r)$ plotted as a function of $T^{*}$ for $\rho^{*}=0.04$, $0.13$, and $0.14$, with an interaction potential characterized by $\epsilon_{1}/\epsilon_{2}=-0.75$, $\lambda_{1}=1.35$, and $\lambda_{2}=1.7$. The circles in panel (a) represent the two states examined in Figure 7 for $\epsilon_{1}/\epsilon_{2}=-0.5$.
@@
Another notable feature is the rounded, bulging profile of the DOC lines for $\epsilon_{1}<0$. This shape indicates that, within a certain density interval, the frequency $\omega$ exhibits reentrant behavior as the temperature varies. This phenomenon is illustrated in Figure 6b for $\epsilon_{1}/\epsilon_{2}=-0.75$. At the density $\rho^{*}=0.04$ (below the loop densities $\rho^{*}\approx 0.05$), the oscillation frequency undergoes a single drop from $\omega\simeq 6.6$ to $\omega\simeq 2.1$ when crossing the temperature $T^{*}\simeq 0.76$ from left to right and then increases smoothly toward $\omega\simeq 4.4$ at the high-temperature limit. At a higher density, $\rho^{*}=0.13$ (just below the bulge’s end at $\rho^{*}\simeq 0.138$), a more complex behavior is observed: the frequency drops from $\omega\simeq 6.6$ to $\omega\simeq 2.5$ at $T^{*}\simeq 1.3$ and then rises again from $\omega\simeq 2.7$ to $\omega\simeq 6.5$ at $T^{*}\simeq 3.6$, eventually tending smoothly toward $\omega\simeq 4.9$ at high temperatures. Finally, at $\rho^{*}=0.14$, the evolution of $\omega$ from $\omega\simeq 6.6$ at low temperatures to $\omega\simeq 4.9$ at high temperatures proceeds continuously without reentrant behavior.
In a manner analogous to the 1D case (see Figure 2), we numerically invert the Laplace transform defined by Equations (13)–(15) using the method outlined in Yuste (2023) to derive $h(r)$. The results for $\epsilon_{1}/\epsilon_{2}=-0.5$, $\lambda_{1}=1.35$, and $\lambda_{2}=1.7$ are presented in Figure 7 for two representative states, labeled A and B in Figure 6a. Additionally, Figure 7 includes the asymptotic form $r|h(r)|=2e^{-\zeta r}|A_{\zeta}\cos(\omega r+\delta)|$ with the parameters $(\zeta,\omega)=(2.324,2.364)$ for state A and $(\zeta,\omega)=(1.283,6.714)$ for state B. The competition between the leading and subleading poles is evident in Figure 7a, where the leading-pole asymptotic behavior requires distances greater than $r\approx 10$. In contrast, for state B, the asymptotic behavior is effectively reached beyond $r\approx 3$. Overall, the contrast between low- and high-frequency oscillations is clearly observed to the left and right of the DOC line, respectively.
The phenomenon of $\omega$ being distinctly larger than $2\pi$ within the loops persists in 3D fluids. However, we have verified that, because these loops are much smaller in size compared to the 1D case, the competition between the leading and subleading poles causes the asymptotic one-pole behavior to dominate only at very large distances. At such scales, the amplitude of the oscillations of $|h(r)|$ can diminish to extremely small values, potentially below $10^{-10}$.
Figure 7: A logarithmic plot of $r|h(r)|$ in the 3D case $\epsilon_{1}/\epsilon_{2}=-0.5$, $\lambda_{1}=1.35$, and $\lambda_{2}=1.7$, for the following states: (a) $(\rho^{*},T^{*})=(0.05,0.6)$ and (b) $(\rho^{*},T^{*})=(0.25,0.6)$. These states are labeled A and B in Figure 6a, respectively. The solid lines illustrate the values derived from numerical Laplace inversion, whereas the dashed lines depict the asymptotic expression $r|h(r)|=2e^{-\zeta r}|A_{\zeta}\cos(\omega r+\delta)|$, where $s=-\zeta\pm\imath\omega$ denotes the leading pole of $G(s)$.
3.4 $\lambda_{1}=1.35$, $\lambda_{2}=1.7$, and $\epsilon_{1}<0$: Influence of $\epsilon_{1}/\epsilon_{2}$ on FW Line
@@
In the 1D case, an FW line is already observed with $\epsilon_{1}/\epsilon_{2}=-0.5$, but this requires temperatures in the order of $T^{*}\sim 10^{-1}$ [see Figure 4a]. Since, as mentioned earlier, the RFA tends to provide less reliable results at low temperatures, it becomes necessary to consider deeper wells to study the FW lines for 3D fluids. The cases $\epsilon_{1}/\epsilon_{2}=-4$, $-8$, $-20$, and $-50$ are reported in Figure 8a. As in the 1D fluid, it is useful to plot the curves on the $T_{1}^{*}$ vs. $\rho^{*}$ plane to compare them with the FW line of the pure square-well fluid, as shown in Figure 8b. Again, we observe that the presence of the repulsive barrier between $\lambda_{1}$ and $\lambda_{2}$ bends the FW line downward for the low-density region. This indicates that the decay of $h(r)$ is always oscillatory when the temperature exceeds a certain threshold, regardless of the density.
@@
Figure 8: (a) FW lines on the $T^{*}$ vs. $\rho^{*}$ plane for the 3D case $\lambda_{1}=1.35$ and $\lambda_{2}=1.7$ with, from the bottom to top, $\epsilon_{1}/\epsilon_{2}=-4$, $-8$, $-20$, and $-50$. (b) The same as in panel (a), except that now the vertical axis represents the scaled temperature $T_{1}^{*}=k_{B}T/|\epsilon_{1}|=T^{*}\epsilon_{2}/|\epsilon_{1}|$. The dotted line is the FW line for a pure square-well fluid ($\epsilon_{2}/|\epsilon_{1}|=0$).
@@ 4 Conclusions
In this paper, we explore the impact of competing interactions in the intermolecular potential of fluids on their structural transitions. The model potential adopted for both 1D and 3D systems consists of a hard core followed by two steps, which can represent either a shoulder or a well depending on the sign of the parameters $\epsilon_{1}$ and $\epsilon_{2}$. This potential is versatile enough to encompass a range of competing interactions, including the square-well and square-shoulder interactions as limiting cases. Additionally, the consideration of two different dimensionalities allows us to examine the influence of strong confinement on the structural transitions of these fluids. For the 1D systems, restricting the interaction range to no more than twice the hard core diameter enables us to derive exact results. In contrast, for the 3D systems, where exact solutions are not feasible, we employ the RFA to obtain and analyze approximate structural properties.
@@
The results for both the 1D and 3D systems align with the expected behavior. Specifically, at very low temperatures, the decay of the total correlation function $h(r)$ exhibits oscillations with a wavelength determined by the range of the repulsive barrier, provided that both $\epsilon_{1}$ and $\epsilon_{2}$ are positive. In contrast, at very high temperatures, the oscillations have a wavelength on the order of the hard core diameter. Furthermore, it is confirmed that at a given density, the transition between these two regimes as the temperature increases can occur either continuously or, as observed in binary hard-sphere mixtures, discontinuously upon crossing a DOC line.
@@
When $\epsilon_{1}$ is negative, an FW transition from an oscillatory to a monotonic decay of $h(r)$ occurs as the temperature decreases at a given density, even when $\epsilon_{2}$ is positive. Additionally, the presence of the repulsive barrier of the height $\epsilon_{2}$ causes the FW line to exhibit a maximum at a certain density before bending downward at lower densities, in stark contrast to its behavior in the absence of such a barrier.
@@
While the results for both the 1D and 3D systems exhibit many common characteristic features, the effects of dimensionality introduce notable distinctions. These include shifts in the temperature ranges in which certain features appear, the need for deeper wells to observe similar phenomena, or a reduction in their prominence as the system transitions from 1D to 3D. In some cases, features present in 1D may vanish entirely in 3D. Notably, we emphasize the complex behavior of the DOC transition, as previously discussed. This intricacy manifests in phenomena such as loops, arcs, lobes, triple points, and reentrant frequencies, some of which, to the best of our knowledge, have not been reported in this context before.
@@
In summary, we uncovered a remarkably complex pattern of structural transitions in fluids with intermolecular potentials that include competing interactions. Even for the relatively simple potential considered in this work, analyzing structural transitions required exploring a broad (dimensionless) parameter space, involving $\lambda_{1}$, $\lambda_{2}$, $\epsilon_{1}/\epsilon_{2}$, $T^{*}$, and $\rho^{*}$. Given these circumstances, our findings are undoubtedly limited. Nevertheless, they reveal a fascinating and intricate phenomenology that merits further and more detailed exploration. In particular, establishing a connection between this phenomenology and the structures and patterns observed in SALR fluids remains an open and compelling challenge.
@@
Conceptualization, A.S.; methodology, S.B.Y., A.S. and M.L.H.; software, A.M.M. and S.B.Y.; validation, A.M.M., S.B.Y., A.S. and M.L.H.; formal analysis, A.S. and M.L.H.; investigation, A.M.M., S.B.Y., A.S. and M.L.H.; writing—original draft preparation, A.S. and M.L.H.; writing—review and editing, A.M.M., S.B.Y., A.S. and M.L.H.; visualization, A.M.M.; supervision, S.B.Y., A.S. and M.L.H.; funding acquisition, S.B.Y. and A.S. All authors have read and agreed to the published version of the manuscript.
@@
This research was funded by MCIN/AEI/10.13039/501100011033, grant number PID2020-112936GB-I00. A.M.M. is grateful to the Spanish Ministerio de Ciencia e Innovación for a predoctoral fellowship, grant no. PRE2021-097702.
[…]
Not applicable.
@@
SALR Short-range attraction and long-range repulsion
@@
Appendix A Some Mathematical Details in the Case of the 1D Fluid